OpenStudy (anonymous):

Simplify. 8xy + 12x – 6x + 13xy

Nnesha (nnesha):

combine like terms

Nnesha (nnesha):

add 0r subtract coefficient of those number who have same variable and same exponent

OpenStudy (anonymous):

so 8xy+13xy-6x+12x?

Nnesha (nnesha):

so combine them add coefficient of xy and add or subtract coefficient of x

OpenStudy (anonymous):

21-18?

OpenStudy (anonymous):

\[21 x y +6 x \]

Nnesha (nnesha):

8 + 13 = ?? 12 -6 = ???

Nnesha (nnesha):

8 + 13 = 21 x right 12 - 6 = x??